OpenStudy (anonymous):

A consumer electronics company sells 6.7 million MP3 players per month at a price of $100 apiece; when the price is lowered to $80, the company sells 10.2 million players per month. How many MP3 players can the company expect to sell per month at a price of $75?

OpenStudy (anonymous):

Let D be the demand and P the price. Then D(100) = 6.7 and D(80) = 10.2. The slope is (10.2-6.7)/(80-100) = -0.175. So using the point-slope formula, the line is D(P) = -0.175 (P-80) + 10.2. When P=75, the demand is D(75) = -0.175 (-5) + 10.2 = 11.075 million.

OpenStudy (anonymous):

thank you :}

OpenStudy (anonymous):

You're welcome :)